OpenStudy (anonymous):

i love brownie !!!!! if i eat 1/3 of brownie today and then 1/4 the left brownie and then 1/9 of the rest ..... how much brownie will be left ????????

OpenStudy (mimi_x3):

LOL, you're obsess of brownies xDD

OpenStudy (anonymous):

i thought it was 1/54

OpenStudy (anonymous):

wow ...... so Good ....Brownies ..

OpenStudy (mimi_x3):

YUMMMM~! xDD *drools*

OpenStudy (anonymous):

8/27

OpenStudy (anonymous):

so many different answers...

OpenStudy (anonymous):

Correction: It is 33/108

OpenStudy (anonymous):

i confirm it...

OpenStudy (anonymous):

I think that is not right: First you eat 1/3, so 2/3 remains Than you eat 2/3*1/4=1/6 2/3-1/6=1/2 Than you eat 1/2*1/9=1/18 1/2-1/18=4/9

OpenStudy (mimi_x3):

LOL, there's no point of solving this stupid question hehe

OpenStudy (anonymous):

right, i agree

OpenStudy (anonymous):

Why? Everyone likes brownies and maths

OpenStudy (mimi_x3):

Not everyone likes maths xD